OpenStudy (anonymous):

I just need help checking my work please...this is a really big grade and I am not confident enough to submit my material...Please Help. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/6 n=1 4*6 = (4(4(1)+1)(8(1)+7))/6 (4(5)(15))/6 =50 24≠50 the base situation doesn’t work, and thus it is proven false.

OpenStudy (anonymous):

2. 12 + 42 + 72 + ... + (3n - 2)2 = (n(6n^2-3n-1))/2 n=1 12= ("1(" 〖6(1)〗^2-3(1)-1))/2 =(6-3-1)/2 =2/2 =1 thus 12=1 so it works for the basic situation n=1 n=k 12+42+72+…+(3k-2)2=(k(k^2-3k-1))/2 assume this is true. n=k+1 12+42+72+…+(3k-2)2+(3k+1-2)2=((k+1)[6(k+1)^2-3(k+1)-1])/2 (k(6k^2-3k-1))/2 +(3k-2)2=((k+1)[6(k+1)^2-3(k+1)-1])/2 (6k^3-3k^2-k)/2 + ((6k-4)^2)/2 = ((k+1)[6(k+1)^2-3(k+1)-1])/2 6k3+15k2+11k+2=(k+1)(6k2+9k+2) I then used synthetic division to factor our k+1 and the results were equal. Thus the statement is proved.

OpenStudy (anonymous):

3. For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1) P1=2=1(1+1)=2 Pk=2+4+6+…+2k=k(k+1) Pk+1=2+4+6+…+2k+2(k+1)=(k+1)(k+1+1)